### v1

# Anything Goes? A Crosslinguistic Study of (Im)possible Language Learning in LMs

Abstract
Do LLMs offer insights into human language learning? A common argument against this idea is that because their architecture and training paradigm are so vastly different from humans, LLMs can learn arbitrary inputs as easily as natural languages. In this paper, we test this claim by training LMs to model impossible and typologically unattested languages. Unlike previous work, which has focused exclusively on English, we conduct experiments on 12 natural languages from 4 language families. Our results show that while GPT-2 small can primarily distinguish attested languages from their impossible counterparts, it does not achieve perfect separation between all the attested languages and all the impossible ones. We further test whether GPT-2 small distinguishes typologically attested from unattested languages with different NP orders by manipulating word order based on Greenberg’s Universal 20. We find that the model’s perplexity scores do not distinguish attested vs. unattested word orders, as long as the unattested variants maintain constituency structure. These findings suggest that language models exhibit some human-like inductive biases, though these biases are weaker than those found in human learners.
Anything Goes? A Crosslinguistic Study of (Im)possible Language Learning in LMs
Xiulin Yang1  Tatsuya Aoyama1  Yuekun Yao2  Ethan Wilcox1 1Georgetown University  2Saarland University {xy236, ta571, ethan.wilcox}@georgetown.eduykyao@coli.uni-saarland.de
1 Introduction
To what extent can LLMs serve as models of human language acquisition and processing? Some, such as Piantadosi (2023), argue that LLMs can function as comprehensive linguistic theories, challenging traditional symbolic generative approaches. On the other hand, critics maintain that the success of LLMs is largely irrelevant to human cognition due to fundamental differences in architecture and learning mechanisms (Chomsky et al., 2023; Fox and Katzir, 2024). Moreover, studies have shown that LLMs fail to acquire key aspects of linguistic knowledge, highlighting their limitations as models of human language (Fox and Katzir, 2024; Lan et al., 2024; Katzir, 2023; Dentella et al., 2024). A particularly compelling argument in this debate is that LLMs are highly flexible learners, capable of acquiring linguistic patterns beyond those learnable by humans, thus making the ability of LLMs to learn human languages uninformative for understanding human language acquisition (Chomsky and Moro, 2022; Moro, 2023; Moro et al., 2023).
We present data that favors a more moderate stance, in line with other researchers in this field (Futrell and Mahowald, 2025; Millière, 2024; Pater, 2019). We present new empirical evidence from the study of impossible languages (Kallini et al., 2024) in a multilingual setting. Our findings suggest that LLMs exhibit learning biases that align with certain aspects of human cognition while simultaneously displaying biases (or a lack thereof) that diverge from human language processing.
In this work, we examine both possible (attested or unattested) and impossible (unattested by definition) languages. Specifically, we define attested languages as the natural languages spoken by humans (e.g., English, German, and Chinese); unattested languages as languages constructed on language universals identified in typological studies as never-occurring; and impossible languages as those that humans cannot acquire and would never produce. Following Kallini et al. (2024), we have selected impossible variants because we take them to be uncontroversial examples of linguistic impossibility, such as languages with randomly shuffled word orders. To explore unattested languages, we draw from Greenberg’s Universal 20 (Greenberg et al., 1963), which identifies unattested word order patterns in noun phrases (e.g., adjective-number-determiner-noun). While there is no direct evidence that such languages are unlearnable, previous studies suggest that typological feature frequencies correlate with learnability in human learners (Culbertson et al., 2020; Gentner and Bowerman, 2009; Saffran et al., 2008).
Regarding impossible language modeling, Kallini et al. (2024) provided initial evidence that GPT-2 small can distinguish between possible and impossible variants of English, suggesting that transformer models encode human-like linguistic biases (Futrell and Mahowald, 2025). However, their study was limited to English, leaving the question of whether this finding generalizes across languages. Furthermore, their focus on impossible languages leaves the study of unattested languages, which we take as an important testbed for LLM-human bias alignment, largely unexplored.
This paper is organized around two main research questions: (1) Do LLMs encode human-like distinctions between attested and impossible languages? Specifically, (a) Within each attested language, can LLMs correctly differentiate the attested language from its impossible variants? (b) Across different attested languages from multiple language families, can LLMs distinguish all attested languages from all impossible languages? (2) Can LLMs recognize unattested languages as distinct from attested ones? Specifically, does LLMs’ ability to model unattested languages align with human typological biases?
Our findings reveal that GPT-2 small reliably distinguishes attested and impossible languages within each attested language (1a) but struggles to make this distinction across different languages (1b). It assigns lower perplexity to unattested languages when they preserve constituency and fixed word order (2), suggesting a human-like preference for regular structures (Hudson Kam and Newport, 2005; Singleton and Newport, 2004) yet deviating from the observed typological biases (e.g., Culbertson et al., 2020). [^1] Our code and data are available at https://github.com/xiulinyang/multilingual-LM.
2 Related Work
2.1 Language Models & Cognitive Plausibility
The advancement of neural networks makes connectionism a widely adopted framework in cognitive language studies (e.g., Wilcox et al., 2023; Borenstein et al., 2024; Kirov and Cotterell, 2018). However, linguists remain divided on whether language models can meaningfully inform linguistic theories. On the one hand, language models have advanced psycholinguistics by serving as highly accurate probability estimators, and, in this capacity, have already been used for testing and refining Surprisal Theory (Goodkind and Bicknell, 2018; Oh and Schuler, 2023b, a; Kuribayashi et al., 2024), Uniform Information Density (Meister et al., 2021; Tsipidi et al., 2024), and other cognitive-linguistic theories and psychometrics (Pearl and Mis, 2011; Gibson et al., 2019; Kuribayashi et al., 2025). On the other hand, their limitations, including a lack of generalization (Yao and Koller, 2022; Kim and Linzen, 2020), the shortcomings of prompt-based approaches (Hu and Levy, 2023), and inconsistency with humans (de Dios-Flores et al., 2023; Davis and van Schijndel, 2020) suggest that, beyond their role as sophisticated estimators, they are limited as cognitive models.
The most relevant work to our study in this context is Kallini et al. (2024), which tests the hypothesis that LLMs cannot distinguish between possible and impossible languages (Chomsky et al., 2023; Moro et al., 2023). Their study relies on a 100M-word dataset from the BabyLM Challenge (Warstadt et al., 2023), focusing on systematically modified versions of English to investigate learnability and model performance. Using the language modeling task with English on GPT-2 small architecture and its impossible variants, Kallini et al. (2024) demonstrate that natural English is consistently easier to learn than its impossible counterparts, as reflected in lower perplexity scores on heldout data. They conclude that the above critique of language models as cognitive models is largely invalid.
2.2 Multilingual Language Modeling
Whether languages vary in complexity remains a controversial topic, and linguists have taken different approaches to address this question (e.g., McWhorter, 2001, 2011; Newmeyer, 2021; Joseph and Newmeyer, 2012). While most generative linguists argue that Universal Grammar requires that all languages be equally complex, others have challenged this notion (Gil, 2008). [^2] See Newmeyer (2021) for a more thorough discussion.
Initial computational attempts to examine language complexity using language models were limited to RNN-based architectures (Cotterell et al., 2018; Mielke et al., 2019; Johnson et al., 2021) and $n$-grams (Koplenig and Wolfer, 2023). These studies suggest that language complexity correlates with morphological richness and the size of speaker populations. More recently, Arnett and Bergen (2025) investigated why morphologically rich languages are harder to model. By testing monolingual language models trained on carefully curated comparative datasets (Chang et al., 2024), they found that morphological features alone could not predict language learnability when training data size was controlled.
While valuable, previous studies often rely on comparative corpora, introducing inconsistencies across languages. Even with parallel corpora (Mielke et al., 2019), studies are limited by small datasets and outdated models. Our study addresses these gaps using a larger parallel corpus and modern transformer architectures.
3 Data and Implementation Details
3.1 Parallel Data Construction: OPUS12 and OPUS30
A key challenge in addressing our questions is that different languages texts drawn from different sources will have different amounts of information. To control for this, we construct two sentence-aligned multilingual parallel corpora to ensure that all languages in our dataset match in terms of content. This allows us to isolate the effect of how formal properties of a language might impact its learnability.
We name the two parallel corpora OPUS12 and OPUS30, gathering aligned sentences from five corpora available on OPUS (Tiedemann, 2012): NLLB (Schwenk et al., 2021), TED2020 (Reimers and Gurevych, 2020), the Bible (Christodouloupoulos and Steedman, 2015), OpenSubtitles (Lison and Tiedemann, 2016), and CCAligned (El-Kishky et al., 2020). Since overlap among languages decreases as more languages are included, we decide to select a minimum of 10M words in English as a standard for our parallel corpora. 10M words also correspond to the amount of input of children’s first 2 to 5 years of development (Warstadt et al., 2023).
OPUS12 is a 12-language multilingual sentence-aligned corpus [^3] The languages and their typological information are listed in Appendix C.. There are around 10M words in the case of English. OPUS30 contains 30 languages with a much smaller size: 48K sentences with 0.7M words. While the two datasets share overlapping languages, their sentences do not overlap, making OPUS30 a suitable test set for additional language modeling experiments.
After deduplicating and removing English sentences from non-English data split using FastText (Joulin et al., 2017), we report the statistics of our corpora in Table 1.
Data Source OPUS12 OPUS30
# Sent # Word # Sent # Word
NLLB 5K 0.1M 16 368
TED2020 164K 2.9M 11K 182K
Bible 40K 1M 14K 324K
OpenSubtitles 680K 4.5M 15K 60K
CCAligned 117K 1.6M 8K  [^1] K
Overall 1M 10.1M 48K 0.7M
Table 1: Data sources of OPUS12 and OPUS30. The word counts are based on the English data. See Appendix C for licensing information.
3.2 Validation Experiment
[…]
3.3 Model Architecture & Training
Group Language Definition
Ours shuffle_local (w=2) The sentence is reordered with every two tokens reversed in order.
reverse_full Every word is reversed in order in a sentence.
k+ shuffle_deterministic (s=84) The sentence is deterministically shuffled by length with seed 84.
shuffle_deterministic (s=57) The sentence is deterministically shuffled by length with seed 57.
shuffle_deterministic (s=21) The sentence is deterministically shuffled by length with seed 21.
shuffle_local (w=10) The sentence is deterministically shuffled in local window size being 10.
shuffle_local (w=5) The sentence is deterministically shuffled in local window size being 5.
shuffle_local (w=3) The sentence is deterministically shuffled in local window size being 3.
shuffle_even_odd The sentence is reordered with even-indexed tokens first, then odd-indexed.
Table 2: Overview of impossible languages in our Experiment1 and Experiment2. k+ languages are borrowed from Kallini et al. (2024) and the rest are new variants introduced in our experiments.
In our experiments, following Kallini et al. (2024), we trained standard GPT-2 small models for each language and evaluated its performance based on the perplexity over a parallel test split of 10K randomly sampled sentences. Due to limited computational resources, we trained each model using 3 random seeds instead of the 5 used in the original study, reduced the maximum training steps from 2000 to 1200 to avoid overfitting, and adjusted the warmup steps proportionally to 120.444We did not experiment with alternative warmup steps, as Kallini et al. (2024) demonstrated that changing the warmup schedule does not affect the ranking of perplexities for impossible language models.
3.4 Multilingual Tokenization
Given our multilingual experiments, tokenization is crucial for fair comparison. To avoid bias toward Latin-script languages, which are overrepresented in our study, we opted against using a multilingual tokenizer with a shared vocabulary.
Previous monolingual experiments either set the vocabulary size of tokenizers to be the same across languages Arnett and Bergen (2025) or applied the formula $0.4\times|V|$ Koplenig et al. (2023); Mielke et al. (2019), where $|V|$ represents the number of unique word types. We conducted a series of pilot experiments on tokenization and found neither approach suitable for our experimental design (Details can be found in Appendix B).
Given these challenges, we decided to use pretrained tokenizers. The rationale behind this choice is that when the tokenizer training data is sufficiently large and diverse, the resulting tokenization scheme should be equally good across languages, as long as the tokenizer algorithm and hyperparameters (e.g., vocabulary size, subword strategy) remain the same. [^5] Although tokenization quality, measured by metrics like compression (Schmidt et al., 2024) and Rényi entropy (Zouhar et al., 2023), has been linked to language modeling performance (e.g., Liang et al., 2023; Goldman et al., 2024), recent studies challenge this connection (Arnett and Bergen, 2025). Some may argue that the BPE algorithm might not be optimized for agglutinative languages such as Turkish, which makes the cross-linguistic comparison unfair. However, much literature on cross-linguistic LM comparison adopts BPE tokenizers (e.g., Mielke et al., 2019; Arnett and Bergen, 2025). As an additional check, we use token counts per word (TCW; reported in Appendix  E Table 5) to measure the morphological complexity of a language and report the correlation between TCW and our test-set perplexity. The results show the correlation is not significant (see Section5), suggesting that the morphological complexity of a language does not substantially impact its learnability.
While it is difficult to say how sufficiently large and diverse a tokenizer training set should be for fair comparison, we consider the size of the training data for GPT2 (Radford et al., 2019) as a reference point, as English was a high-resource language even in 2019 when the paper was published. We believe that this data size is sufficient to minimize differences that tokenization will make across languages.
When selecting pretrained tokenizers, we primarily use monolingual BPE tokenizers. [^6] However, for Chinese, we follow previous studies (Mielke et al., 2019) and use the Chinese-BERT tokenizer. Our goal is to maintain a relatively consistent vocabulary size across languages (50k), though we make exceptions for Romanian, Arabic, and Chinese due to limited model availability. The training data for all other languages is at least as large as the English corpus. The tokenizer details can be found in Appendix D.
Figure 1: Attested individual Language vs. their corresponding counterparts with a 95% confidence interval over 3 random seeds tested on 10k sentences from OPUS30.
4 Experiment 1: Attested vs. Impossible Languages in an Intra-Language Modeling Setting
4.1 Impossible Languages
In this experiment, we use the deterministic shuffled languages from Kallini et al. (2024) along with two new variants (see Table 2). We include shuffled languages because (1) Kallini et al. (2024) identify them as the most impossible languages in their language possibility ranking, and (2) their difficulty is also indirectly supported by empirical studies showing that both adults and children exhibit a regularization bias, which can be thought of as a bias against shuffling (Newmeyer, 2005; Singleton and Newport, 2004).
Since all languages are deterministically shuffled, the original ones (i.e., attested ones) can be recovered from their variants through another deterministic function. If LLMs function as non-human-like pattern recognizers as Chomsky et al. (2023); Moro et al. (2023) argue, they should be able to learn these deterministic languages.
4.2 Results & Discussion
Based on the results presented in Figure 1, we have three key observations. First, in all languages except Italian, the perplexity of the attested language is lower than all its impossible variants. For Italian, shuffle_local (w=2) yields a slightly lower perplexity than natural Italian, though the difference is not statistically significant (Mann-Whitney U test: W = 63, p = 0.353). Additionally, attested languages exhibit smaller error bars compared to their impossible counterparts, indicating more stable learning. Third, we observe a consistent learnability pattern across languages. For each language, smaller shuffling windows result in lower perplexity. Moreover, most languages perturbed with shuffle_deterministic are harder to model than shuffle_local ones. We assume this is because shuffle_deterministic languages are shuffled based on sequence length and autoregressive models do not have direct access to this information when predicting the next word, making language modeling harder.
We also conducted a Spearman’s Ranking Correlation test between the results on OPUS30 English and those from Kallini et al. (2024)’s experiments. We grouped the shuffle_deterministic languages together and observed that the ranking of our English impossible variants aligns perfectly with that reported by Kallini et al. (2024) ($\rho=1$, $p=0.0027$).
Based on these findings, we answer the first subquestion: language models can (largely) distinguish between each attested language and their corresponding impossible counterparts.
5 Experiment 2: Attested vs. Impossible Languages in an Inter-Language Modeling Setting
In this experiment, we gather the learning results of all possible and impossible languages to see if there is a separation boundary between them. If GPT-2 small can distinguish between possible and impossible languages, we expect the former’s perplexity to be lower than the latter’s.
Langs Attested Example
Typo. Theo.
perturb_Nnda no no She enjoyed books three the fantastically interesting  a lot .
perturb_anNd no no She enjoyed fantastically interesting  three books the a lot .
perturb_daNn few yes She enjoyed the fantastically interesting  books three a lot .
dperturb_dnaN many yes She enjoyed the three fantastically interesting  books a lot .
perturb_dnNa many yes She enjoyed the books three fantastically interesting  a lot .
np_random no no She enjoyed books fantastically three interesting  the a lot .
Table 3: List of NP-perturbations with corresponding categories and examples. Typo refers to typologically-attested, while Theo refers to theoretically-attested.
Figure 2: Attested natural languages vs. impossible languages with a 95% confidence interval over 3 random seeds. The x-axis represents the training steps, and the y-axis shows the perplexity on the test split. All the impossible languages are marked in light blue.
The results of different language models are shown in Figure 2.777To highlight the overlap of perplexity between attested languages and impossible ones, we zoom in on the lower perplexity range while displaying higher perplexity values in a separate, compressed section with a break in the $y$-axis. According to the figure, we see that not every language is equally easy to learn: Chinese is the easiest language, while Arabic is the hardest, followed by Turkish and Russian. We observe a moderate positive correlation between the average number of tokens per word (TCW) and perplexity of each of the last checkpoints in 11 languages (Chinese is excluded because the BERT tokenizer is a character-level tokenizer), as indicated by a Spearman’s rank test ($\rho=0.564$), but it is not significant ($p=0.076$). This finding aligns with the observation by Arnett and Bergen (2025) that there is no significant difference in language modeling difficulty of agglutinative vs. fusional languages when the amount of information is controlled. Although all the attested languages are distributed at the bottom of the graph, we see some impossible languages fall between these attested languages. For example, Russian, Turkish and Arabic all show higher perplexity than English perturbed with shuffle_local (w=3). This means that for GPT-2 small, these impossible languages are just as hard to learn as the attested languages.
To quantify the extent GPT-2 small can distinguish attested from impossible languages, we train a linear SVM classifier with the perplexity value across the three random seeds of each checkpoint as features. The classifier reaches $0.75$ ($sd=0.08$) macro F1 score averaged over 10-folds cross-validation.
Based on this experiment, we answer the second sub-question posed in our paper: Although language models tend to learn attested languages better than impossible ones, their perplexity does not distinguish all attested languages from all impossible languages overall.
Figure 3: Attested individual natural languages vs. their corresponding impossible languages with a 95% confidence interval over 3 random seeds. The x-axis represents the training steps, and the y-axis shows the perplexity on the test split. Different language types are distinguished using distinct color palettes.
6 Experiment 3: Attested vs. Unattested Languages
In this experiment, we investigate how well language models can learn unattested languages, languages whose structure is conceivable according to rules of grammar or morphology, but which have not been attested. While unattested languages are not necessarily unlearnable (e.g., Tsimpli and Smith, 1995), prior research suggests a link between typological feature frequency, cognitive biases, and language learnability (e.g., Gentner and Bowerman, 2009; Culbertson et al., 2012; Culbertson and Newport, 2015; Culbertson et al., 2020).
We focus on Greenberg’s Universal 20 (Greenberg et al., 1963), which suggests that certain determiner-adjective-number-noun orders in an NP are universally unattested. As a well-studied typological phenomenon, Universal 20 serves as a good testbed for comparing human learners and language models, complementing experimental findings that show harmonic NP orders (i.e., ones where the dependents always all either precede or follow the head; e.g., num-adj-noun and noun-adj-num) are easier to learn than non-harmonic ones (e.g., num-noun-adj or adj-noun-num) (Culbertson and Newport, 2015, 2017; Culbertson et al., 2020). One influential hypothesis, the Typological Prevalence Hypothesis, proposes that more common typological patterns are easier to learn (Gentner and Bowerman, 2009). Therefore, we predict that if language models exhibit similar biases as humans, we expect a gradient of difficulty in learning different NP orders, with some unattested configurations posing greater challenges than others.
Among the 24 theoretically possible orders of adjectives, nouns, determiners, and numbers, we select five combinations, covering cases classified as few, many, and zero in Cinque (2005)’s typological analysis. [^8] Although Cinque (2005) seeks to explain why zero languages really are “underivable” under the minimalist program we refer to them as unattested to contrast them with the impossible languages of the previous section, i.e., ones that involve shuffling or reversed word order. In this experiment, we only permute words within NPs. If the perplexity of these permuted languages is similar to that of attested languages, it suggests two possible reasons: (1) Language models can learn these unattested languages; (2) the number of words in NP may be a small number with respect to the entire data size, and hence NP-internal perturbation introduces a much smaller noise compared to the entire data perturbation we used in previous experiments, which may not significantly affect the learnability of a language. To rule out the latter possibility, we also construct a control condition in which words corresponding to these POS categories are randomly shuffled within NPs. This language serves as a baseline, indicating the extent to which NP-internal permutations influence the learnability of a language.
Examples of perturbed NP word orders and their typological information are listed in Table 3 and their word orders are reported below:
  • perturb_Nnda: noun>num>det>adj.
  • perturb_anNd: adj>num>noun>det.
  • perturb_daNn: det>adj>noun>num.
  • perturb_dnaN: det>num>adj>noun, typical of English and Chinese.
  • perturb_dnNa: det>num>noun>adj, typical of Italian and Portuguese.
  • np_random: Random permutation of adj, noun, num, and det within NPs.
Since identifying NP structures requires a constituency parser, we use Stanza (Qi et al., 2020) to parse raw text. Stanza provides constituency parsing for only Chinese, Portuguese, English, and Italian, with acceptable accuracy (>0.85) [^9] https://stanfordnlp.github.io/stanza/constituency.html, so we limit our analysis to these four languages. As different parsers are trained on distinct treebanks with varying annotation guidelines, we select POS tags based on each treebank’s guidelines. Details are provided in Appendix F.
Results
Our results are visualized in Figure 3 (bottom subgraph). Surprisingly, all five NP-perturbed languages exhibit lower perplexity compared to their attested counterparts across all four languages. Two of these (Nnda and anNd) are unattested in typological studies and are ruled out by generative approaches (Cinque, 2005), but we do not observe a significant difference in perplexity between the three languages with attested NP orders and the two languages with unattested orders.
When these POS tags are shuffled within NPs, perplexity increases, reaching or exceeding the perplexity of the attested languages. This rules out the possibility that limited perturbations simply do not affect model training.
To summarize, this experiment shows that, unlike humans, language models fail to show a gradient of difficulty in learning different NP orders according to their typological prevalence.
Discussion: Why Can’t LMs Distinguish Between Attested and Unattested Languages?
To make sense of these results, we propose two key factors that influence LM learning outcomes: randomness and constituency structure. By randomness, we refer to whether the perturbation function produces a perturbed text that can be deterministically recovered to its original form. By constituency structure, we mean whether the phrase structures of the original language are preserved in the perturbed version.
Regarding randomness, string distributions with higher entropy are always harder to learn. This explains why NP-perturbed unattested languages show lower perplexity than attested languages and np_random variants. The reasoning is that our perturbation procedure enforces a strict ordering procedure, which may be (sometimes) violated in the original attested language. For example, although English is a dnaN language, certain constructions such as the dann (det-adj-num-noun; e.g., a beautiful five days) does not follow the dominant pattern. Once POS tag orders are normalized within NPs, the resulting constructions become more predictable. Therefore, all normalized NPs, including our unattested NPs, may have lower overall entropy, which could explain why they are easier to learn. In fact, the normalized dnaN, which has the same typical word order as English, shows lower perplexity than the original, unnormalized English; and the same applies to our other languages in this experiment.
Regarding constituency structure, we hypothesize that disrupting constituency weakens local dependency relations within phrase structures. This explains why all impossible languages in the previous experiments, despite maintaining a deterministic order, still results in higher perplexity than NP-perturbed languages (Figure 3). Similarly, this may also explain the higher perplexity of count-based grammars in Kallini et al. (2024): the insertion of morphological markers disrupts phrase structure integrity. One exception is reverse_full, which preserves constituency structure while maintaining a deterministic order. As shown in Figure 3, reverse_full exhibits perplexity closer to unattested languages. We hypothesize that this exception may be due to other factors, such as information density (Clark et al., 2023). Since reversing word order alters information flow, it may obscure more accessible information. For example, predicting a pronoun given a preceding noun is easier than predicting a noun given a preceding pronoun, potentially increasing difficulty for the LM. One shortcoming of this experiment is that it was based entirely on NP perturbations(Kallini et al., 2024). Future work could investigate whether similar effects occur when POS order or constituency structure is disrupted on a larger scale.
7 Conclusion
In this paper, we extend Kallini et al. (2024) to a broader multilingual context. Our experiments provide mixed results that complement those of Kallini et al. (2024). We find that while GPT-2 small can distinguish between attested languages and their impossible variants, its learning outcomes do not separate all attested and all unattested or impossible languages. That being said, LMs do tend to learn attested languages better, on average, than impossible languages, and we achieve a separability of $0.75$ between the two classes based on the models’ perplexity. Finally, we observe that some unattested languages show lower perplexity than their attested counterparts even though they exhibit NP orderings that flout Greenberg’s Universal 20.
What to make of these results in the context of our original question–whether LLMs can serve as cognitive models? While our results show that GPT-2 does not behave as we might expect from a fully human-like learner, they also demonstrate that it has a soft preference for attested over impossible languages. Skeptics have previously likened LLMs to a bad theory of physics in which “anything goes.” [^10] Chomsky, quoted from an email to Gary Marcus: You can’t go to a physics conference and say: I’ve got a great theory. It accounts for everything and is so simple it can be captured in two words: “Anything goes.” All known and unknown laws of nature are accommodated, no failures. Of course, everything impossible is accommodated also. In line with Kallini et al. (2024), our results demonstrate that these models do not instantiate an “anything goes” hypothesis. Rather, their incremental data-processing architectures represent a useful starting point for studying human language processing. Refining models to achieve stronger alignment with people is possible, and will likely lead to lasting insights about human cognitive architecture.
8 Limitations
We acknowledge that our experiments rely on GPT-2 Small, which may not generalize to larger models. This choice was made for two reasons: (1) running experiments across multiple languages is computationally expensive; (2) we aimed for comparability with Kallini et al. (2024). Future work could explore whether our findings hold for larger models. Additionally, the dataset used for training the language model is relatively small. This is a deliberate trade-off between data size and linguistic diversity. While a larger dataset might yield more robust results, our approach ensures broader typological coverage. Lastly, in our experiments on unattested languages, we generated synthetic data by perturbing languages based on Universal 20. However, linguistic correlations extend beyond word order universals. For instance, Greenbergian correlations (Dryer, 1992) suggest that verb-object order often correlates with other features such as adposition-noun phrase order and determiner-noun phrase order. Future work will explore more nuanced perturbations to better capture such cross-linguistic dependencies.
9 Ethics Statement
We use publicly available datasets, ensuring that no private or personally identifiable information is included. Our dataset selection prioritizes linguistic diversity while maintaining data transparency. Regarding computational resources, we use GPT-2 small trained on A-100 and V-100 GPUs. Each experiment on each language took around 10-12 hours.
Appendix A Experiment Results of Replicating Kallini et al. (2024)
We implement the training and evaluation following the same experiment setting from Kallini et al. (2024), only on a 10M word subset of their original data. The result is shown in Figure 4. Unlike in Kallini et al. (2024), however, we do observe that test-set perplexity does increase towards the end of training, indicating that models are overfitting on our smaller datasets. We note that we do not observe this overfitting behavior in the experiments presented in the main text, where the heldout perplexity continues to decrease (or plateau) throughout training.
We calculate Spearman’s rank correlation between our results for the *shuffled languages and those of Kallini et al. (2024) at every 200-step interval from 400 to 1,200. The Spearman’s $\rho$ is consistently 1 (p < 0.001), indicating perfect agreement between the rankings, showing that 10M words are sufficient enough to replicate the language modeling experiments for which Kallini et al. (2024) originally used 100M words.
Figure 4: Replication of (Kallini et al., 2024) with 10M words from BabyLM Challenge dataset (strict-small track)
Figure 5: Perplexity results on the development set (10K sentences) for five languages (EN, TR, RO, DE, RU), trained on a 10M-sentence training set across different vocabulary sizes. Error bars represent the first and last quartiles (25% and 75%) of the results. A plot for the optimized vocabulary size (labeled ‘BEST’) is also included, showing high variance for TR and RU even with optimized vocabulary size.
Language Treebank POS-tags
det num adj noun
English Penn Treebank (Marcus et al., 1993) dt, prp\$, pdt, pos qp, \$, cd rb, adjp, jjr, jjs, jj nn, nns, nnp, nnps
Italian VIT(Delmonte et al., 2007) det num, sq adj, sa noun, pron, propn, sym, x
Chinese CTB 3.0(Xue et al., 2005) dt, m, clp, dp cd, od, qp jj, adjp, dnp, dec, deg nn, np, nr, nt, prp, pn, fw
Portuguese Cintil (Barreto et al., 2006) det, d, dem, poss, poss’ qnt, qnt’, num, percentp, percentp’, card, card’ adj, ap n’, noun, pron
Table 4: POS-tag categories across languages
langs ar tr ru pl de it
tcw 2.19 2.05 2.05 1.98 1.65 1.40
langs pt nl ro en fr
tcw 1.68 1.51 1.81 1.45 1.67
Table 5: TCW per language by each of their pretrained tokenizer
Appendix B Tokenization Pilot Experiments and Results
In our experiments, where we trained tokenizers for each language using 10M words (around 60MB data), testing vocabulary sizes ranging from 30K to 80K in increments of 10K, we observed two key findings: (1) Tokenizers trained with around 60MB data resulted in unstable language modeling outcomes, and (2) different languages require distinct optimal vocabulary sizes: morphologically richer languages tend to have a larger vocabulary size. We also observed that even when trained on the corpus with matching content, not all languages are equally learnable in terms of their perplexity. These results are shown in Figure 5. Additionally, agglutinative languages like Turkish, with their large number of unique tokens, made large vocabulary sizes impractical. For instance, Turkish has three times the number of unique words as English (467K vs. 140K), and applying $0.4\times|V|$ would result in a vocabulary size of 186K, which is too large for efficient language model training with the limited data available and a small model.
Language Family Word Order Morphology
OPUS12
English Indo-European (Germanic) SVO Analytic
German Indo-European (Germanic) No dominant Fusional
Russian Indo-European (Slavonic) SVO Fusional
Romanian Indo-European (Romance) SVO Fusional
Turkish Turkic (Altaic) SOV Agglutinative
Dutch Indo-European (Germanic) No dominant Fusional
Polish Indo-European (Slavonic) SVO Fusional
Portuguese Indo-European (Romance) SVO Fusional
Italian Indo-European (Romance) SVO Fusional
French Indo-European (Romance) SVO Fusional
Chinese Sino-Tibetan SVO Analytic
Arabic Afro-Asiatic (Semitic) VSO Root-based (nonconcatenative)
OPUS30
Spanish Indo-European (Romance) SVO Fusional
Czech Indo-European (Slavonic) SVO Fusional
Bulgarian Indo-European (Slavonic) SVO Fusional
Slovak Indo-European (Slavonic) SVO Fusional
Serbian Indo-European (Slavonic) SVO Fusional
Croatian Indo-European (Slavonic) SVO Fusional
Ukrainian Indo-European (Slavonic) SVO Fusional
Danish Indo-European (Germanic) SVO Fusional
Swedish Indo-European (Germanic) SVO Fusional
Greek Indo-European (Hellenic) No dominant Fusional
Persian Indo-European (Indo-Iranian) SVO Fusional
Lithuanian Indo-European (Baltic) SVO Fusional
Vietnamese Austroasiatic SVO Analytic
Hebrew Afro-Asiatic (Semitic) VSO Root-based (nonconcatenative)
Hungarian Uralic SVO Agglutinative
Indonesian Austronesian SVO Analytic
Japanese Japonic SOV Agglutinative
Korean Koreanic SOV Agglutinative
Table 6: Typological features of the OPUS12 and OPUS30 corpora, with OPUS30 including 18 additional languages beyond those in OPUS12.
Appendix C Details of OPUS12 and OPUS30
The typological features of languages used in the two corpora are reported in Table 6. The licensing terms vary depending on their original sources, listed below.
Appendix D Tokenizers
Table 7 shows the details of the tokenizers we use in the experiments. When the training data for a tokenizer is unspecified, we assume it matches the training data used for the corresponding pretrained model.
Appendix E TCW
The TCW is reported in Table 5. We use it to measure the morphological richness of a language.
Appendix F POS tags of each treebank
Different constituency parsers are trained with different treebanks. We select POS-tags that are relevant to the four word classes. The detailed POS-tags for each language can be found in Table 4.
Language |Vocab| |Training| Reference Domain
Arabic [^1] https://huggingface.co/aubmindlab/aragpt2-base 64,000 77GB Antoun et al. (2021) Web Crawl, Wikipedia, News
Turkish [^2] https://huggingface.co/ytu-ce-cosmos/turkish-gpt2 50,257 100GB Kesgin et al. (2024) Web Crawl, books, news, others
Russian [^3] https://huggingface.co/ai-forever/rugpt3large_based_on_gpt2 50,257 450GB Zmitrovich et al. (2024) Wikipedia, books, news, books, Web Crawl, Subtitles
Polish [^4] https://huggingface.co/flax-community/papuGaPT2 50,257 47GB Wojczulis and Kłeczek (2021) Web Crawl
German [^5] https://huggingface.co/malteos/gpt2-xl-wechsel-german 50,304 156GB Ostendorff (2023) Web Crawl
Italian [^6] https://huggingface.co/iGeniusAI/Italia-9B-Instruct-v0.1 50,176 Trillions toks iGeniusAI (2024) public sources, synthetic data, and domain-specific content
Portugese [^7] https://huggingface.co/NOVA-vision-language/GlorIA-1.3B 50,258 35B tokens Lopes et al. (2024) Web Crawl, News, Subtitles, EuroParl
Dutch [^8] https://huggingface.co/yhavinga/gpt-neo-125M-dutch 50,257 151GB Havinga (2023) Web Crawl
Romanian [^9] https://huggingface.co/dumitrescustefan/gpt-neo-romanian-780m 64,000 40GB Dumitrescu (2024) Web Crawl, Opus, Wikipedia
English [^10] https://huggingface.co/openai-community/gpt2 50,257 40GB Radford et al. (2019) Web Crawl
French [^11] https://huggingface.co/lightonai/pagnol-xl 50,262 130GB Launay et al. (2022) Web Crawl
Chinese [^12] https://huggingface.co/google-bert/bert-base-chinese 21,128 300GB Devlin et al. (2019) Wikipedia
Table 7: Tokenizers, vocabulary sizes, training data sizes, references, pretrained model name, and training data domains for each language tested in our experiments.